A freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hour slower than the express train. They pass one another 3 hour later. What are their speeds?
Thanks for your help.

Answer by venugopalramana(3286)

Initial distance between 2 trains=390 km
Final distance between 2 trains=0 km….(since they pass one another)
Hence distance traveled by both together=390-0=390 km
Let speed of express train =x kmph
30 kmph slower than above=x-30 kmph=speed of goods train
since both are traveling in opposite directions ,their relative speed is sum of their speeds =x+x-30=2x-30 kmph (note that with each passing hour,when the trains move in opposite directions,their distance of separation comes down by he sum of their speeds)
time of travel =3 hrs.
so we have…..relative speed =distance trvelled between final and initial stgs/time difference from beginning to end of travel
2x-30=390/3=130
2x=130+30=160
x=160/2=80
speed of express = 80 kmph
speed of goods = x-30 =80-30=50 kmph